if three forces f1 =20n, 30 degree north east, f2 = 50n along west, f3 =40n, 50 degree north west act on a body. find the resultant in magnitude and direction


1
Expert's answer
2021-03-16T11:36:25-0400

Let's find the "x"- and "y"-components of the resultant force:


"F_{res,x}=F_{1x}+F_{2x}+F_{3x},"

"F_{res,x}=20\ N\cdot cos30^{\circ}+50\ N\cdot cos180^{\circ}+40\ N\cdot cos(180^{\circ}-50^{\circ})=-58.4\ N,"


"F_{res,y}=F_{1y}+F_{2y}+F_{3y},"

"F_{res,y}=20\ N\cdot sin30^{\circ}+50\ N\cdot sin180^{\circ}+40\ N\cdot sin(180^{\circ}-50^{\circ})=40.64\ N."

We can find the magnitude of the resultant force from the Pythagorean theorem:


"F_{res}=\sqrt{F_{res,x}^2+F_{res,y}^2},""F_{res}=\sqrt{(-58.4\ N)^2+(40.64\ N)^2}=71.15\ N."

We can find the direction of the resultant force from the geometry:


"\theta=sin^{-1}(\dfrac{F_{res,y}}{F_{res}}),""\theta=sin^{-1}(\dfrac{40.64\ N}{71.15\ N})=34.8^{\circ}."

The direction of the resultant force is "34.8^{\circ}\ N\ of\ W."